OpenStudy (anonymous):

Solve x2 + 4x + 8 = 0.

OpenStudy (anonymous):

x equals negative 2 plus or minus 2 I x equals negative 2 plus or minus 4 I x equals negative 4 plus or minus 2 I x equals negative 4 plus or minus 4 I

OpenStudy (anonymous):

lets get this right, make sure you understand go step by step \[x^2+4x+8=0\] step one subtract \(8\) from both sides what do you get ?

OpenStudy (imstuck):

@satellite73 don't you think that the quadratic formula would be easier than completing the square for someone who is not proficient at factoring?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x2+4x=0 ? @satellite73

OpenStudy (imstuck):

ahh...difference of opinion then. My students in Algebra 2 hated completing the square once they knew the quadratic formula!

OpenStudy (anonymous):

a) the middle term is \(4x\) and 4 is even there will be no denominator the quadratic formula forces a denominator on you then you have a raft more algebra to do

OpenStudy (anonymous):

b) most people can't remember the formula correctly

OpenStudy (anonymous):

c) almost no one can do the arithmetic right

OpenStudy (anonymous):

x equals negative 2 plus or minus 2 I x equals negative 2 plus or minus 4 I x equals negative 4 plus or minus 2 I x equals negative 4 plus or minus 4 I

OpenStudy (imstuck):

but understanding competing the square is so much more difficult for students. And who's to say they will remember completing the square? Especially when you factor out a number to make the coefficient on the x^2 a 1.

OpenStudy (anonymous):

@brangetsbuckets no it is not \[x^2+4x=0\] if you start with \[x^2+4x+8=0\] the \(8\) cannot disappear subtract \(8\) from both sides

OpenStudy (imstuck):

But you're a brillliant mind and I bow to your brains. I can see it both ways personallly; I enjoy completing the square so much more, but my students hate it.

OpenStudy (anonymous):

@IMStuck thanks for the compliment i guess i should have said in a case like this it is much easier to complete the square if you had \(7x^2+5x-2=0\) the no it would not be

OpenStudy (imstuck):

you are welcome...I meant it with total sincerity. Go on with your helping; I'll butt out now! ; )

OpenStudy (anonymous):

im still lost @satellite73